OpenStudy (ams98):

I need help, I worked out a problem but I'm not sure of how to display the answer! I've never seen it this way! HELP! MEDAL?!

OpenStudy (ams98):

\[(\sqrt[4]{x^2})(\sqrt[4]{x^2}) \] \[(x\frac{ 2 }{ 4 })(x\frac{ 2 }{ 4 }) \] \[x \frac{ 4 }{ 4}\] My question is how to put this as a radical...?

OpenStudy (ams98):

I know that mean x plus one whole other x so is that just (x) or (x^2) or radical X..???

OpenStudy (anonymous):

so you have eventually x^(4/4) = x^1 = x

OpenStudy (anonymous):

so as i said your found that the answer is x^(4/4) it is the same as x^(1) which is the same as x